OpenStudy (anonymous):

what is the standard form of the equation of the circle with center (3,-2) and a point on the circle (-1,1)

hartnn (hartnn):

can u find the distance between those 2 points ? that would be the radius r

OpenStudy (anonymous):

using the distance formula

hartnn (hartnn):

yes

OpenStudy (anonymous):

i got 13

hartnn (hartnn):

how?

hartnn (hartnn):

3--1 = 4 , 4^2=16 -2-1=-3, -3^2-9

hartnn (hartnn):

-3^2=9***

OpenStudy (anonymous):

i didnt square the 4

OpenStudy (anonymous):

so its 25

hartnn (hartnn):

sqrt 25 =...?

OpenStudy (anonymous):

5

hartnn (hartnn):

you have centre , u have radius r=5, can you find the equation now?

OpenStudy (anonymous):

is it (x+1)^2+(y-1)^2=5

hartnn (hartnn):

the center is 3,-2 not -1,1

hartnn (hartnn):

\((x-h)^2+(y-k)^2=r^2\)

hartnn (hartnn):

h=3,k=-2

OpenStudy (anonymous):

so its (x-3)^2 and (y+2)^2= 25

hartnn (hartnn):

yes + instead of and

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^